Current assets 1,333,333

Long-term assets 4,666,667

Total assets 6,000,000

Current liabilities 637,754

Long-term liabilities 3,577,724

Owners' equity 1,784,522

Salmon Enterprises

Bonds​ outstanding:  

3,000

selling at

​$999.37

Common stock​ outstanding:  

260,000

selling at

​$30.99

Book value versus market value components.  Compare​ Trout, Inc. with Salmon​ Enterprises, using the balance sheet of Trout and the market data of Salmon for the weights in the weighted average cost of​ capital:   If the​ after-tax cost of debt is 10.3​% for both companies and the cost of equity is 15.55​%, which company has the higher​ WACC? a. What is the book value adjusted WACC for Trout, Inc.?
____% (Round to two decimal places.)

What is the market value adjusted WACC for Salmon​Enterprises? ____%

Which company has the higher​ WACC? A. Salmon Enterprises has a higher WACC at ___ than​ Trout, Inc. with a WACC of ___ B. Trout, Inc. has a higher WACC at ___ than Salmon Enterprises with a WACC of ___
Answer #1

Trout Inc. Component weights Debt = 8,577,724 / 5,362, 246 = 0.667 Equity = 1,784,522 / 5,362,246 = 0.333 Salomon Enterprises